# On the contraction properties of Sinkhorn semigroups

### v1

@@ 3.2 Some contraction coefficients
$\jmath({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~{}\inf_{y\in\mathbb{Y}}~{}\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
$\hbar({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (24)
@@ 3.2.1 Birkhoff contraction coefficient
[…]
As remarked in Section 3 of [22], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t. the Hilbert metric.
3.2.2 Dobrushin contraction coefficients
[…]
$\chi_{\varphi,\psi}({\cal K})=\sup_{\nu_{1},\nu_{2}\in{\cal M}_{1,\varphi}(S)}\frac{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}{\cal K}-\nu_{2}{\cal K}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}-\nu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}}$
@@ 3.2.2 Dobrushin contraction coefficients
[…]
@@ 3.2.3 Entropy contraction coefficient
[…]
For any convex function $\Phi$ satisfying (21) we recall that
[…]
The estimate (32) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [21] for finite state spaces and [37] for general measurable spaces). For further details on the Dobrushin contraction coefficient we refer the reader to the article [41], see also [35, 37] and Chapter 4 in [34].
4 Sinkhorn contraction properties
4.1 Some regularity conditions
@@
$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~{}:~{}\varphi(x)\leq r\right\}.$
When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$. Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [6] and references therein). In the context of Gibbs-loop semigroups (14), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
Condition (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
@@
$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
@@
$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.$ (37)
@@
4.2 Norm contraction theorems
@@
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, then there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that
@@
The contraction estimate (38) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $0\leq p\leq n$ we readily obtain
$\|(\mu_{1}-\mu_{2}){\cal S}^{\circ}_{2p,2n}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\mu_{1}-\mu_{2}\|_{\psi_{\kappa}}.$
Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\pi_{0},\nu_{V})$, the above estimate immediately yields an upper bound on the total variation distance, i.e.,
$\|\pi_{2n}-\nu_{V}\|_{\sf tv}\leq\|\pi_{2n}-\nu_{V}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\pi_{0}-\nu_{V}\|_{\psi_{\kappa}}.$
Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [30, 70, 71]
$\|{\cal S}^{\circ}_{0,2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{0,2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{n}~{}\left(\psi(y_{1})+\psi(y_{2})\right).$
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may make it hard to test –except in the case of Gaussian and linear models. Several Lyapunov functions in this context are designed in Section 5.2. A stronger (but easier to test) condition can be provided in terms of the Sinkhorn transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
Condition (${\cal D}_{\varphi,\psi}({\cal S})$).
@@
$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
@@
When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, we readily see that
${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y})$
for any $n\geq n_{0}$.
@@
If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (33) and (34) also hold for any $n>n_{0}$. In addition, there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that, for any $n\geq n_{0}$,

### v2

@@ 3.2 Some contraction coefficients
$\jmath({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~\inf_{y\in\mathbb{Y}}~\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
$\hbar({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (38)
@@ 3.2.1 Birkhoff contraction coefficient
$\raisebox{2.0pt}{$\chi$}_{H}({\cal K}):=\sup\frac{H(\nu_{1}{\cal K},\nu_{2}{\cal K})}{H(\nu_{1},\nu_{2})}=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}},$ (40)
${\raisebox{2.0pt}{$\chi$}_{H}({\cal K})<1}\Longleftrightarrow\hbar({\cal K})>0\Longleftrightarrow\jmath({\cal K})>0.$
We also emphasize that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}$ there is an interesting twist because the uniform minorization condition $\jmath({\cal K})>0$ is generally not satisfied. We illustrate this assertion with the semi-circle transition on the unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ defined by
${\cal K}(x,dy)=\frac{1}{q(x)}~\sqrt{1-(x-y)^{2}}~1_{]0,1[}(y)~dy$ (41)
for some normalizing constant $1/3\leq q(x)\leq 1$. In this case, choosing $x_{1}=1/(2n)=1-y=1-x_{2}$ in (37) we check that
$0\leq\jmath({\cal K})\leq 3~\sqrt{1-(1-1/n)^{2}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0\Longrightarrow\jmath({\cal K})=0.$
The entropic transport model (4) on the unit interval ]0,1[]0,1[ with reference transition (41) is discussed in Section 5.4. Note that in that case, the cost function is unbounded.
3.2.2 Dobrushin contraction coefficients
Consider functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ for any $(x,y)\in(\mathbb{X}\times\mathbb{Y})$. The $(\varphi,\psi)$-Dobrushin coefficient $\raisebox{2.0pt}{$\chi$}_{\varphi,\psi}({\cal K})$ of a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by the norm operator
@@ 3.2.2 Dobrushin contraction coefficients
The terminology $\varphi$-Dobrushin coefficient comes from the fact that we recover the standard Dobrushin contraction coefficient $\raisebox{2.0pt}{$\chi$}({\cal K})$ by choosing constant functions. Indeed, if $\varphi=1=\psi$, then
$\raisebox{2.0pt}{$\chi$}_{1,1}({\cal K})=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{\|\nu_{1}{\cal K}-\nu_{2}{\cal K}\|_{\sf tv}}{\|\nu_{1}-\nu_{2}\|_{\sf tv}}=:\raisebox{2.0pt}{$\chi$}({\cal K})=1-\varepsilon({\cal K}),$ (43)
@@ 3.2.3 Entropy contraction coefficient
Given a convex function $\Phi$ satisfying (12), the Dobrushin $\Phi$-contraction coefficient $\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})$ of some Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K}):=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{D_{\Phi}(\nu_{1}{\cal K},\nu_{2}{\cal K})}{D_{\Phi}(\nu_{1},\nu_{2})},$ (46)
where the supremum is taken over all probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ such that $D_{\Phi}(\nu_{1},\nu_{2})>0$. Here, again, the terminology $\Phi$-Dobrushin coefficient comes from the fact that choosing $\Phi(u,v)=\Phi_{0}(u,v)=\frac{1}{2}|u-v|$ we recover the standard Dobrushin contraction coefficient, i.e., $\raisebox{2.0pt}{$\chi$}_{\Phi_{0}}({\cal K})=\raisebox{2.0pt}{$\chi$}({\cal K})$.
For any convex function $\Phi$ satisfying (12) we recall that
$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})\leq\raisebox{2.0pt}{$\chi$}({\cal K}).$ (47)
The estimate (47) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [cohen-0] for finite state spaces and [dm-03] for general measurable spaces). For further details on the Dobrushin contraction coefficient we refer the reader to the article [dob-56], see also [dg-01, dm-03] and Chapter 4 in [dm-04].
4 Sinkhorn contraction properties
4.1 Strong mixing models
As underlined in [fienberg, mosteller] in the context of finite state space models, Sinkhorn iterations of the form (16) preserve the cross product Radon-Nikodym ratio in (38) and, hence, for any $n\geq 0$ we have
$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}).$ (48)
This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [borwein, chen, deligiannidis, franklin]. Recall that the Hilbert contraction coefficient of a Markov transition $\cal Q$} can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq. (40)), hence the exponential convergence of Sinkhorn iterations w.r.t. the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$. An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
From (39), let us observe that
Thus, for any given convex function $\Phi$ satisfying (12), we can use (40) and (47) to show that
$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
while
@@ 4.1 Strong mixing models
Theorem 4.1.
The above inequalities remain valid when we replace $(\nu_{V},\pi_{2n})$ by $(\lambda_{U},\pi_{2n+1})$. If we assume that $\jmath({\cal Q})>0$, then $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ and, in addition, there exists some $c>0$ such that
Proof.
$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
Corollary 4.2.
Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [dg-01]. For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself. The choice ${\cal Q}(x,dy)~\propto~e^{-\delta\,\|y-b(x)\|}~dx$ yields
In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [aronson], Nash [nash] and Varopoulos [varopoulos] on Gaussian estimates for heat kernels on manifolds. Nevertheless, the above condition is not met for Gaussian transitions.
4.2 Gibbs-Loop Sinkhorn semigroups
@@
$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~:~\varphi(x)\leq r\right\}.$ (52)
When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$.
Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [dpa] and references therein). In the context of Gibbs-loop semigroups (23), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
Condition 4.3 (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
@@
[…]
@@
$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.\$(57)
@@
4.3 Weighted norm contractions
@@
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied some integer $n_{0}\geq 1$, then there exist parameters $\kappa>0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (53) and on the function $\varepsilon(r)$ in (54) and such that
@@
The contraction estimate (58) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $n_{0}\leq p\leq n$ we readily obtain
[…]
Note that
[…]
Choosing $p=n_{0}$ and $(\mu_{1},\mu_{2})=(\pi_{2(n_{0}-1)},\nu_{V})$, the above estimate immediately yields, for any $n\geq n_{0}$, an upper bound on the $\psi$-norm. We summarize the above discussion with the following corollary.
Corollary 4.5.
Under the assumptions of Theorem 4.4 for any $n\geq n_{0}$ we have
[…]
with the parameters $\rho\in[0,1[$ and $\kappa>0$ as in (58).
Choosing $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [persi, saloff-zuniga, saloff-zuniga-2]
$\|{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{1+n-n_{0}}~\left(\psi_{\kappa}(y_{1})+\psi_{\kappa}(y_{2})\right).$
Remark 4.6.
The proximal sampler discussed in Remark 2.1 has a Markov transition ${\cal S}^{\circ}_{1}$ with invariant measure $\lambda_{U}$. The weighted-norm contraction properties of these models are simply based on the existence of a Lyapunov function $\varphi$ satisfying the r.h.s. condition in (53) with $n=0$. As in (54) we also require a local minorization condition $\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{1}\|_{\sf tv}\leq 1-\varepsilon(r)$ on the sub-level sets $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$ of the Lyapunov function to ensure that $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{1})<1$, for some parameter $\kappa\geq 0$ (see for instance Theorem C.2).
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may not be easy to test – except in the case of Gaussian and linear models. Several Lyapunov functions in this context are designed in Section 5.8. A stronger (but easier to test) condition can be provided in terms of the transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
Condition 4.7 (${\cal D}_{\varphi,\psi}({\cal S})$).
@@
[…]
@@
When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, for any $n\geq n_{0}$ we readily see that
${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y}).$
In addition, we have
$\pi_{2n+1}(\varphi)=\nu_{V}{\cal S}_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c$ (62)
[…]
@@
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds as soon as condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied and the estimates (59) hold for some $n_{0}\geq 1$ and some parameters $\rho\in[0,1[$ and $\kappa>0$.
When ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied there also exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (60) and on the function $\varepsilon(r)$ in (61) and such that, for any $n\geq n_{0}$,
$\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa},\psi_{\kappa}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa},\varphi_{\kappa}}({\cal S}_{2n+1})<\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
A proof of Theorem 4.8 is provided in Appendix C (on page C).
Let us emphasize that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is stronger than Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$. In particular, the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ can be stable even when the Sinkhorn transitions ${\cal S}_{n}$ are unstable. Intuitively, Gibbs-loop transitions go forward and backward towards their invariant measures (22) while the Sinkhorn transitions in (16) transport asymptotically a target marginal measure to the other. For instance, by Lemma 4.9 in [adm-24] (see also Section A) the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ of the Gaussian Sinkhorn algorithm are always stable. However, in the context of regularized Gaussian entropic problems on a time horizon $t$ (cf. Section 3.4 in [adm-24]) the Sinkhorn transitions ${\cal S}_{n}$ may be unstable when the time horizon is not sufficiently large (by Remarks 3.9 and 4.6 in [adm-24]).
4.4 A Lyapunov stability theorem
Regularity conditions
This section presents some sufficient conditions for ${\cal D}_{\varphi,\psi}({\cal S})$, in terms of the potential functions $(U,V,W)$. These simple conditions enable us to obtain geometric rates for the convergence of the Sinkhorn iterations for a variety of models, including polynomial growth models (Section 5.1) and heavy tailed models (Section 5.2), as well as boundary states models including Beta and Weibull distributions (Section 5.3 and Section 5.4). Illustrations in the context of Linear-Gaussian models and finite (non necessarily Gaussian) mixture marginals are also discussed respectively in Section 5.8 and Section 5.9.
Let us assume that $U,V,W$ are locally bounded and $e^{-W}\in\mathbb{L}_{\infty}(\lambda\otimes\nu)$. In addition, $U,V$ have compact sub-level sets. These conditions ensure there exists some $W_{\star}\in\mathbb{R}$ such that
$W(x,y)\geq W_{\star}\quad\mbox{\rm for $(\lambda\otimes\nu)$-almost every $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.}$ (63)
In addition, $(U,V)$ are bounded below by a real number so that $e^{-U}$ and $e^{-V}$ are uniformly bounded. This excludes probability densities that tends to infinity at boundary states, such that some classes of Weibull or the Beta densities for some shape parameters. Also note that for any $\delta>0$ we have
$(e^{\delta U},e^{\delta V})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
Consider the following conditions:
Condition 4.9 (${\cal H}_{\delta}(U,V)$).
There is some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ defined in (27) are locally bounded with compact sub-level sets and we have
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$ (64)
Note that (63) and (64) together imply the inequalities
$\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)\leq e^{-W_{\star}}~\left(\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\right)<\infty.$
This yields the exponential moment estimates
${\cal H}_{\delta}(U,V)\Longrightarrow\lambda_{(1-\delta)U}\left(1\right)=\lambda_{U}\left(\exp{(\delta U)}\right)<\infty\quad\mbox{\rm and}\quad\nu_{(1-\delta)V}(1)=\nu_{V}\left(\exp{(\delta V)}\right).$
In particular all absolute moments $\lambda_{U}(|U|^{p})\vee\nu_{V}(|V|^{p})<\infty$, with $p\geq 0$ are finite. Given the reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ also have
${\cal H}_{\delta}(U,V)\Longrightarrow\mbox{\rm Ent}((\lambda_{U}\otimes\nu_{V})~|~{\cal P}_{0})=(\lambda_{U}\otimes\nu_{V})(W)-\nu_{V}(V)<\infty.$ (65)
To check this claim, note that
$(\lambda_{U}\otimes\nu_{V})(W)$ $=$ $\lambda_{U}(W^{V})=W_{\star}+\lambda_{U}(W^{V}-W_{\star})$
$\leq$ $W_{\star}+\lambda_{U}(e^{W^{V}-W_{\star}})=W_{\star}+e^{-W_{\star}}~\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}).$
On the other hand, for any $r>0$ we have
$\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}})$ $\leq$ $e^{-r}~\lambda_{(1-\delta)U}(1)+\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}~1_{{\cal U}_{\delta}\leq r})$